### v1

# Independent researcher, Grenoble, France

11institutetext: Univ. Lille, Inria, CNRS, Centrale Lille, UMR 9189 CRIStAL, F-59000 Lille, France 22institutetext:
Independent researcher, Grenoble, France
33institutetext:
LIP6, Sorbonne Université, F-75005, Paris, France
Easing Maintenance of Academic Static Analyzers
Raphaël Monat    Abdelraouf Ouadjaout    Antoine Miné
(Submitted: 17/07/2024)
Zusammenfassung
Academic research in static analysis produces software implementations. These implementations are time-consuming to develop and some need to be maintained in order to enable building further research upon the implementation. While necessary, these processes can be quickly challenging. This article documents the tools and techniques we have come up with to simplify the maintenance of Mopsa since 2017. Mopsa is a static analysis platform that aims at being sound. First, we describe an automated way to measure precision that does not require any baseline of true bugs obtained by manually inspecting the results. Further, it improves transparency of the analysis, and helps discovering regressions during continuous integration. Second, we have taken inspiration from standard tools observing the concrete execution of a program to design custom tools observing the abstract execution of the analyzed program itself, such as abstract debuggers and profilers. Finally, we report on some cases of automated testcase reduction.
Keywords:
Static Analysis Abstract Interpretation Software Engineering
journal: Journal
1 Introduction
One of the products of academic research in static analysis is the implementation of tools, which are used to illustrate and evaluate approaches. During their lifetime, these tools have to be maintained – in particular to enable further research. This research can be performed by the same authors, new members of a same group or by other groups wanting to build upon this tool. While necessary, debugging and maintenance of static analyzers can quickly turn out to be time-consuming, as static analyzers perform highly technical reasoning. In addition, this maintenance is not the main purpose of academic jobs – and similarly, industrial developers will be looking to reduce their maintenance costs as much as possible. Due to its nature, the maintenance problem is common to all researchers in the community. Some practices are considered as folklore, which might explain why practices developed after years of experience are not systematically documented nor shared between groups. In this paper, we document the tools and techniques we use to simplify our maintenance of Mopsa, the static analysis platform we have been developing since 2017. We hope this article, following the precursor work of Andreasen et al. (2017), will have a twofold impact, by being useful to other researchers (and especially newcomers such as students), and by motivating other groups to share their practices.
Contributions.
  • We describe a novel way of reporting analyses results, and measuring precision based on selectivity (Section 3). It is automatic, enhances transparency of the analysis, and takes into account the complexity of the expressions in the analyzed program (and not merely the program size). We can then leverage analyses results to detect soundness and precision regressions, thanks to a set of real-world, open-source benchmarks used in our continuous integration.
  • In Section 4, we show how plug-in observers to the analysis can help in the development of coverage and profiling tools working on the abstract interpretation of the program, while standard tools provide different results by working on the concrete execution of the analyzer.
  • We showcase an abstract debugger interface allowing interactive exploration of the abstract interpretation of the analyzed program, facilitating use and debugging (Section 5). This abstract debugger is available both as a command-line interface and in IDEs supporting the Debug Adapter Protocol.
  • We report on our experience in applying automated testcase reduction tools (initially used on compilers) on static analyzers (Section 6). These tools ease the process of isolating, analyzing and fixing precision and soundness issues in the analyzer. Indeed, this process is essentially manual and time-consuming, and bugs are often initially identified on example runs too large to inspect manually. Collaboration between Mopsa and automated testcase reduction is a two-way street: Mopsa enables seamless testcase reduction for multi-file projects with complex commands such as GNU coreutils. This process is enabled thanks to the mopsa-build utility instrumenting the compilation process, which can then be leveraged to output a single preprocessed file usable by testcase reduction tools. The generated file corresponds to a kind of AST-level linking.
In particular, Sections 4 and 5 highlight a systematic connection between standard tools observing the concrete execution of the abstract interpreter and custom tools (abstract debuggers, profilers) we developed, which observe the abstract execution of the analyzed program itself.
2 An Overview of Mopsa
Mopsa (Modular Open Platform for Static Analysis) is a publicly-available (Miné et al. (2024)) and open-source framework for the development of static analyzers based on the theory of abstract interpretation by Cousot and Cousot (1977). Journault et al. (2019); Monat (2021) describe its flexible and modular architecture that makes it extensible in many aspects. Mopsa aims at simplifying the exploration of new ideas and the development of static analyzers, while providing mature implementations for selected mainstream languages. Mopsa participated in 2023 and 2024 to the Software-Verification Competition (SV-Comp), following contributions from Monat et al. (2023, 2024b). In 2024, Mopsa obtained the first place in the SoftwareSystems track of SV-Comp.
2.1 Languages
Mopsa supports the analysis of multiple languages, such as C (Journault et al. (2018); Ouadjaout and Miné (2020)), Python (Monat et al. (2020a, b)), combination of Python and C (Monat et al. (2021)) and Michelson (Bau et al. (2022)) languages. Unlike most multi-language static analyzers, Mopsa does not rely on translating programs to a fixed intermediate representation before starting the analysis. These intermediate representations are generally very low-level (e.g., three-address code, stack machine, small C subsets). While intermediate representations simplify the reuse of abstractions among different languages, they may suffer from information loss – such as high-level control-flow structures and data structures – during the initial syntactic translation. For example, LLVM forgets whether integer types are signed or unsigned, while transformation to 3-address code puts a strain on relational domains to maintain precision, as shown by Namjoshi and Pavlinovic (2018). In addition, a fixed intermediate representation may not support all kinds of languages and paradigms (e.g, dynamically typed, object oriented programming languages, functional languages), limiting the generality of the framework.
On the contrary, Mopsa has an extensible AST (Abstract Syntax Tree) that represents the union of the original ASTs of the supported languages. This approach preserves the original semantics of the program, allowing developers to reason on it directly. However, Mopsa is not a disjoint union of separate analyzers, but tries to reuse abstractions between different languages. This is done via two key mechanisms: semantic rewriting and delegation.
Instead of an initial syntactic translation, as performed by classic analyzer frontends, Mopsa can rewrite a statement to another languages during the analysis. This dynamic translation can exploit the inferred values of variables to produce more precise, or more efficient transformations. For example, an integer addition in C can be translated to a mathematical addition if the sum of the operands fits within the range of the expression type. Similarly, an addition in Python can be translated to a mathematical addition if the operands have only values of integer types, and do not implement the special methods __add__ and __radd__. Mopsa provides ready-to-use abstractions for mathematical integers, such as intervals and polyhedra. Therefore, both C and Python analyses can delegate the processing of the translated expression to one of these domains, which enables abstraction reuse between different analyses.
2.2 Precision
Sound static analyzers may be imprecise due to coarse over-approximations. This problem can be overcome by designing more accurate abstractions. These improvements are generally very localized, targeting specific language constructs, such as how integer values are encoded or how loops are iterated. Therefore, it is important to decompose the global abstraction into smaller pieces, called domains, that are responsible for different parts of the language. Domains handling the same statements or expressions should be easily exchangeable without impacting the remaining analysis, or seamlessly combined in a reduced product to enhance precision.
In Mopsa, the definition of the global abstraction is done with a JSON file called a configuration, that lists the domains to include in the analysis as well as their order of execution and relations; and domains implement a common unified OCaml signature.
Unified Signature.
From a developer perspective, the unified signature simplifies the integration of new domains. This signature provides an API to allow cooperation between domains while preserving low coupling among them. For example, when delegating the execution of a statement, a domain cannot call another domain directly. Instead, it asks the framework to find within the configuration the appropriate domain(s) implementing the transfer function of the statement.
Configurations.
From a user perspective, the JSON format of configurations gives a simplified way to define a new analysis by combining existing domains. Mopsa supports different types of combinators to compose domains, such as sequences and reduced products. Many ready-to-use configurations with various tradeoffs between precision and efficiency are also provided.
2.3 Properties
Similarly to most static analyzers, Mopsa can verify classic reachability properties, such as absence of runtime errors in C and uncaught exceptions in Python. In addition, some analyses are experimented in Mopsa, for in-progress research projects. In particular, some target more complex reachability properties. Parolini and Miné (2024) verify the user-exploitability of alarms in C codes. Delmas et al. (2021) developed ways to prove portability of C programs between architectures with different endianness. Monat et al. (2024a) have developed an analysis targeting the rounding-sensitivity of date computations, in the context of legal implementations using the Catala programming language (Merigoux et al. (2021)).
These kinds of reachability properties rely on computing necessary post-conditions with an over-approximating forward analysis. The aim of these kinds of analysis is certifying the correctness of the program. Mopsa was recently extended to support the computation of sufficient preconditions via an under-approximating backward analysis described by Milanese and Miné (2024). This kind of analysis is able to generate counter-examples for certifying program incorrectness.
3 Measuring Precision
The precision of static analyzers is a cornerstone of experimental evaluations developed to evaluate the benefits of new approaches. It can also be leveraged to detect changes and regressions during tool development. We provide a quick survey of classic approaches to measuring precision in static analysis, before introducing a new way to automatically compute precision, which improves the transparency of the analysis and which can be numerically quantified. Then, we highlight how our approach can be naturally leveraged to compare analysis results, and how this comparison is used to detect changes during development, through continuous integration.
3.1 Traditional Approaches to Measuring Precision
The precision of a static analyzer is usually judged by separating the true bugs it found from the number of false alarms it raised. [^1] Note that in some cases, precision is measured through the proxy of another construction, such as generated call graphs Smaragdakis et al. (2011); Helm et al. (2024). In practice, this measure requires a baseline to be established. This baseline requires tedious manual work discriminating alarms, which is almost impossible to realize on new analyses of large-scale projects. We start by describing the case of manually annotated benchmarks where an absolute precision measure can thus be computed. We then survey usual ways to measure precision when no baseline is known – i.e, the number of true bugs is not known.
Absolute precision: the case of manually annotated benchmarks.
Some specific benchmarks have been manually crafted, or studied, to know where the true bugs lie. This is for example the case of NIST’s Juliet test suite for static analyzers by Black (2018), which contains tests labeled either as safe (no runtime errors) or buggy (with a single runtime error). In that case, establishing an absolute precision measure is possible, provided of course that the classification is correct. We leverage parts of the Juliet test suite to detect potential regressions of Mopsa in our continuous integration (cf. Section 3.4). We show in Table 1 our current precision results, measured as the percentage of tests where Mopsa is optimally precise.
When no baseline exists: usual approaches to measuring precision.
In most cases however, manually checking all alarms to separate true alarms from the others is not possible for the purpose of experimental evaluations made in academia. In those cases, other precision measures exist. A first measure, which can be useful on small testcases, is recording a boolean holding when the program is safe. This measure however is extremely coarse, which makes it unpractical to use during experimental evaluations or to discover improvements in an analysis. A second approach is to report the number of alarms. In our experience, this absolute number is however not informative: the numbers are difficult to put in context and might puzzle the community. This can be mitigated by measuring number of alarms per total lines of code, but this measure will be highly application-dependent.
Note that other approaches to quantify precision can be leveraged when the goal is to compare the results of two analyses of a same program. One can measure and compare abstract states at similar points of the analyses. This measure however yields more questions than it solves: which abstract states should be compared? For example, comparing all abstract states in the same contexts might be time-consuming due to the sheer number of states. In addition, some precision changes will propagate from one state to its successor(s), which may skew the measure. Additional issues can arise when incomparable domains are used. In related work, the clam static analyzer from Gurfinkel and Navas (2021); Gurfinkel et al. (2015) provides a clam-diff utility to compare two different analyses of a same program, which relies on a semantic comparison of the numerical abstract states computed in each analysis. This approach has more granularity than the one we currently use, and Arceri et al. (2023) have used it to evaluate precision changes in their work. However, this approach may be too sensitive and it could be interesting to quantify the difference between some comparisons. For example, the abstract state $a_{1}=x<1$ is included in $a_{2}=x\leq 1$, which itself is included in $a_{3}=\top$, but the change from $a_{2}$ to $a_{3}$ denotes a bigger precision loss. To the best of our knowledge, this problem of quantifying the difference between two abstract states has only been considered by Sotin (2010).
3.2 Reporting static analysis results in an automatic and transparent fashion
{minted}
[escapeinside=||]ocaml (* |$a^{\#}$| abstract state, |$p^{\#}$| safety property *) if |$a^{\#}\not\sqsubseteq p^{\#}$| then add_alarm |$a^{\#}$| |$p^{\#}$|
[…]
Abbildung 1: High-level implementation change, to move from reporting alarms to a transparent report of alarms and successful checks.
Abbildung 2: Analysis report summary for the analysis of coreutils fmt.
{minted}
[escapeinside=||,numbers=left,xleftmargin=2em]c int main(int argc, char** argv) int y = -1; for(int x = 0; x < argc; x++) y++;
(a) Toy C example.
Stmt. Analysis Intervals Polyhedra
x++ Safe Safe
y++ Alarm Safe
Selectivity 50% 100%
(b) Selectivity measurement, in the case of integer overflow detection for the toy example of Figure 3(a), analyzed either using intervals or a relational polyhedra abstract domain.
Abbildung 3: Illustrating selectivity computation on a toy C example.
We have developed an approach to report static analysis results in Mopsa which is both automatic and enhances the transparency of the verifications performed by the analysis.
Traditionally, static analyzers only report alarms, which correspond to failed proofs of safety: program locations222and callstacks, for our fully context-sensitive analyses. where the abstract state does not satisfy a property the analyzer checks (such as absence of runtime errors). Our approach consists in also logging successful proofs, which we name safe checks (of a given a property). In the implementation, this change is conceptually easy to add, and boils down to the snippets shown in Figure 1. The analysis can then report all checks it performed, including alarms, and in a complementary way, the number of safe checks. An example of real analysis report is shown in Figure 2. From this report, we can derive a numeric notion of precision we call selectivity, measuring the ratio of successful proofs Mopsa has been able to perform. In the case of the analysis of coreutils fmt in Figure 1, the selectivity is 87%.
Selectivity=#checks proved $\text{Selectivity}=\frac{\#\text{checks proved safe}}{\#\text{checks}}$
We illustrate these notions of (safe) checks and selectivity computation on a toy example in Figure 3. In that case, we assume we are only interested in checking integer overflows, so we only check the two increments x++ and y++. In a configuration using a non-relational interval analysis, Mopsa will be able to prove that x stays in the range of signed integers, but will fail to prove it for y. In that case, the selectivity is thus one half. Moving to a relational analysis relying on polyhedra, Mopsa will be able to infer that y is bounded by x, and hence prove that both operations are safe from overflows, resulting in a selectivity of 100%.
This approach has several benefits: it requires a lightweight implementation and provides transparent results, where users can clearly read in the analysis reports what the analysis has been able to verify. Finally, selectivity provides a relative measure of precision, making it more informative than reporting an absolute number of alarms. This measure depends on the complexity of the analyzed program expression and statements, rather than more arbitrary measures such as a program size.
3.3 Comparing analyses results using mopsa-diff
[…]
[fontsize=]textmopsadiff_touch.txt
(a) Comparing two analyses on a single program.
[…]
[fontsize=]textmopsadiff_multi.txt
(b) Comparing two analyses on all coreutils.
Abbildung 4: Mopsa-diff output comparing the impact of different relational domains on some coreutils programs.
mopsa-diff is a tool we developed which can compare analyses reports: it can either focus on comparing two analyses of a given program, or on comparing analyses results on a set of programs. Comparing two analyses of a common program can be used to detect breaking changes (in terms of soundness or precision) when the implementation of the analyzer changes. Similarly, we can compare the performance-precision benefits of changing configurations. mopsa-diff can also be lifted to inspect the impact of a configuration change to a set of benchmarks for example.
These two modes are displayed in Figure 4, where we compare the analysis of some coreutils program with two relational numerical domains (chosen for illustrative purposes): PPLite v0.11 from Becchi and Zaffanella (2020) and the NewPolka implementation of Apron v0.9.14 from Jeannet and Miné (2009). Figure 4(a) compares two analyses of coreutils touch. The diff-like output shows that the PPLite version is able to remove 11 alarms compared to new Polka, although it is a bit slower, and adds another alarm. Figure 4(b) provides an overall comparison on all coreutils. We can notice that a few alarms are removed by this version, although two new alarms are generated. The analysis is overall slower. No new soundness assumptions [^3] During an analysis, Mopsa may make specific assumptions impacting the soundness of the analysis. Mopsa takes great care to report any such assumptions, and issues a special warning for the user to check them when there are some, for the sake of transparency. For example, Mopsa assumes by default that external C functions have no side effects on their parameters or global variables. We believe this approach to be a practical take on the principles highlighted in the soundiness manifesto of Livshits et al. (2015). are recorded, and the PPLite analysis does not crash on any program NewPolka is able to analyze.
3.4 Detecting breaking changes during continuous integration
We leverage the analyses reports and mopsa-diff in our continuous integration to detect breaking changes affecting the soundness and precision of the analysis. This is done by comparing obtained results with baseline results. The set of benchmarks used in our continuous integration corresponds to all open-source projects we have analyzed in past experimental evaluations of our works. We strive not to modify the source code in order to stick close to a real-world usage. In some highly exceptional cases, we may rely on stubs to improve the results. We currently have just one stub for all benchmarks used in our continuous integration. Tables 1, 2 and 3 show the benchmarks we currently use. Running times have been measured on a desktop machine relying on an Intel Core i7-12700.
Benchmark # Tests Total LOC Time Precision
CWE121 2,508 234,930 3,064s 22.13%
CWE122 1,556 166,664 1,948s 25.84%
CWE124 758 93,372 961s 36.94%
CWE126 600 75,984 769s 46.83%
CWE127 758 89,022 963s 37.07%
CWE190 3,420 440,749 4,356s 78.13%
CWE191 2,622 340,884 3,236s 78.87%
CWE369 497 83,238 674s 70.42%
CWE415 190 17,990 228s 100.00%
CWE416 118 14,782 142s 67.80%
CWE469 18 1,520 22s 100.00%
CWE476 216 20,427 254s 100.00%
Tabelle 1: Results of Mopsa analysis on Juliet benchmarks (non-relational configuration, no partitioning). CWE121 contains 2,508 tests, which overall take 3,064s to analyze. Mopsa is able to pass 22.13% with a precise analysis, and is imprecise in the 77.87% remaining cases.
Benchmark LOC (C) LOC (Python)
coreutils 148,214 0
pyperformance 0 4,215
fpp 0 3,140
bitarray 2,969 2,474
cdistance 912 979
levenshtein 5,120 357
llist 2,757 1,686
noise 636 631
pyahocorasick 2,933 1,336
Tabelle 2: Link and lines of code (loc) of each project Mopsa uses in its continuous integration. LOCs have been measured using cloc from Danial (2021).
Benchmark Time Selectivity # checks
coreutils basename 33.79s 98.65% 11,731
comm 42.67s 97.32% 12,654
dircolors 34.82s 99.74% 20,062
dirname 21.68s 99.61% 11,307
echo 19.26s 99.43% 11,010
false 14.50s 99.72% 10,774
getlimits 34.62s 98.54% 11,711
hostid 18.05s 99.65% 11,303
id 32.69s 99.04% 12,338
link 23.03s 99.52% 11,572
logname 20.36s 99.66% 11,307
mkfifo 34.87s 99.20% 11,807
mknod 34.98s 99.11% 12,513
nice 23.36s 99.55% 11,463
nohup 26.98s 99.27% 11,734
nproc 17.37s 99.44% 11,533
printenv 23.59s 99.50% 11,202
pwd 22.04s 99.62% 11,502
rmdir 39.00s 99.22% 11,699
runcon 18.55s 99.66% 11,215
seq 42.68s 95.87% 14,310
sleep 23.79s 99.46% 11,546
stdbuf 32.16s 98.46% 12,526
sync 24.53s 99.60% 11,273
tee 35.69s 98.76% 12,057
timeout 32.28s 98.51% 12,420
true 9.55s 99.72% 10,774
uname 20.61s 99.52% 11,943
unlink 16.17s 99.63% 11,497
users 20.82s 99.06% 11,668
whoami 13.03s 99.66% 11,329
yes 19.82s 99.45% 11,216
pyperformance chaos 8.55s 98.86% 6,415
fannkuch 0.31s 98.75% 321
float 0.10s 99.48% 574
go 143.72s 97.67% 7,552
hexiom 39.30s 98.33% 5,392
nbody 1.28s 99.73% 1,100
raytrace 49.99s 98.95% 5,694
regex_v8 18.30s 99.56% 32,638
richards 11.86s 99.65% 3,710
scimark 22.74s 99.09% 6,397
spectral_norm 1.30s 99.43% 697
unpack_sequence 3.46s 100.00% 8,094
fpp choose 155.27s 99.76% 91,417
processInput 3.84s 99.78% 4,914
bitarray 499.26s 89.57% 288,475
cdistance 58.26s 96.54% 46,512
levenshtein 27.26s 85.35% 15,519
llist 78.70s 98.92% 136,237
noise perlin 5.80s 99.20% 4,273
simplex 6.06s 99.45% 4,586
pyahocorasick 29.92s 89.67% 20,415
Tabelle 3: Benchmarks used in Mopsa’s non-regression testsuite. coreutils benchmarks have been analyzed with fully symbolic arguments, and a relational analysis. Other projects have been analyzed in a non-relational setting.
4 Instrumenting the Analysis
Hooks are plug-ins that can observe, and influence, the analysis. When enabled, these hooks are called before and after every statement is analyzed, and they can peek at the input and output abstract states. Hooks do not have access to the private, internal representation of abstract domains, but communicate through a public interface of queries (described by Journault et al. (2019)).
These hooks have a wide variety of usages ranging from providing interpretation traces to improving the precision by detecting relevant thresholds which can be used by the widening. In the context of debugging and maintaining a static analyzer, these hooks are handy because they work at the level of the abstract execution of the input program, while standard profiling and coverage tools can provide information about the execution of the static analyzer. In the remainder of this section, we showcase computations of abstract coverage, abstract profiling, and heuristic detection of unsoundness and large imprecision.
4.1 Coverage
This hook computes the abstract coverage of statements that have been reached by the analysis. It gives a birds-eye view of where the analysis went, and can easily be leveraged to detect insufficient modeling assumptions reducing the search space of the analysis, as the user can quickly spot functions that should be reachable but are not. These modeling assumptions can be due to the modeling of command-line arguments.
As an example, the analysis of coreutils fmt shows that 76% of its main function is covered, when the analysis simply assumes that no arguments are passed to the utility. Given that fmt is a command-line utility, assuming that no arguments are passed is too restrictive. We can thus move to an analysis which performs a symbolic modeling of arguments: it considers that the analyzed program arguments are an array of arbitrary size (within either system bounds or user-supplied bounds) containing strings of arbitrary size and contents covering all actual possible usage. In this case of symbolic modeling of arguments, we reach 100% coverage for the main function. Note that by default, the analysis makes no restriction on the arguments of the main entry point (i.e, arguments are modeled symbolically by default).
This hook can help users find soundness issues related to their configuration and instrumention of the analysis.
4.2 Profiling
Abbildung 5: Flamegraph obtained using the abstract function profiling of Mopsa, when analyzing coreutils fmt.
Loops and function calls are the two reasons why the same block of code might need to be analyzed a large number of times (due to loop iterations in one case, and context-sensitivity in the other case). They are thus the main source of analysis cost, and it is important to be able to pinpoint which ones might be problematic in a specific analysis. Mopsa thus provides two profiling hooks: one for loops and one for function calls.
The loop hooks tracks the number of times a given loop is called, the number of iterations needed to reach a fixpoint, as well as the total time spent analyzing each loop. This is helpful to fine-tune the widening, to track slowdowns in the postfixpoint computation performed by Mopsa, and more extremely, non-termination issues.
The function profiling hook tracks the number of times a function is analyzed as well as the time spent analyzing it. This hook is particularly important as the current analyses written in Mopsa are fully context-sensitive, meaning they analyze functions by virtual inlining. The hook thus helps identifying which functions take the most time to analyze. In particular, the collected data can be transformed into the flamegraph graphical representation from Gregg (2016); an example is shown in Figure 5. Tracking the number of times the function is analyzed is also important, as a frequent reason a function is reported as taking a large time is when the function is called many times, within nested loops, even if each function analysis is actually fast.
Hooks can produce and return information during abstract computations, without having to wait for the analysis to terminate. In particular, they work on partial executions of the analysis, which is especially useful for profiling programs where the analysis is unexpectedly long. It will only give a partial picture but will highlight the relevant loops and functions taking time to be analyzed.
4.3 Heuristic Unsoundness/Imprecision Detection
We developed plugins performing heuristic detection of unsound or highly imprecise behaviors, which are reported to the user. Most users will not precisely know the behavior of all domains an analysis configuration uses, especially due to the highly distributed nature of transfer functions in Mopsa. It is thus interesting to have hooks acting as runtime mechanisms that can warn the user of unsound or highly imprecise behavior, making it easier to pinpoint a source of imprecision or unsoundness.
The unsoundness detector acts as a sanity check. It currently verifies the following property: an assignment from a non-bottom state cannot return bottom [^4] Except if the expression contains a runtime error, such as a division by zero.. Other properties could easily be added to this detector in future work. While the coverage hook finds modeling and configuration errors from users, this detector finds soundness issues due to bugs in the analyzer. These errors are critical, and corrected as soon as possible, as the analyses are constructed to be sound.
The imprecision detector warns when the analysis of a (sub)expression yields top, which is the source of large imprecision. Precision issues are less critical in Mopsa, as they usually result in a performance-precision tradeoff.
5 Abstract Debugging through an Interactive Engine
Traditional approaches to debugging static analyzers do not scale well. We briefly survey basic techniques – all supported by Mopsa – and their shortcomings. A first approach is to check the analysis output. It is usually quite coarse, and when unexpected behaviors happen, the output is not sufficient to get explanations. Another approach is to provide some builtin functions asking for the abstract interpreter to print its current state. Those can then be added in the source program to show specific abstract states. Bühler et al. (2024) mention the use of Frama_C_show_each in Frama-C; Mopsa provides a similar mopsa_print. While this approach helps making the result understandable, the cost is somewhat high: the modification of the source code requires to restart the analysis each time, which can be prohibitive when programs are large. If the program location where the printing happens is reached in a lot of different contexts, the output may turn out to be too verbose to be useful. Additionally, this restart of the analysis can become more complicated if the precise location of the origin of the error is different from the location where it becomes manifest. In that case, it must be discovered by trial and error, by inserting logging commands and running the analysis again many times. A complementary approach is to store relevant analysis states and information so that users can inspect them afterwards. Thus, a single analysis run can produce a log that can be examined at different program locations, even if these locations are not known in advance. The Goblint tool from Saan et al. (2024a) provides an interactive output allowing to explore various abstract states after the analysis is finished, either as an HTML page or through the Debug Adapter Protocol available in different IDEs and pioneered by VSCode. In Mopsa, we can record an interpretation trace, showing the order in which expressions and statements are analyzed, optionally with the abstract state. This however can quickly become too big to process. For example, the interpretation trace of analyzing coreutils fmt in Mopsa is around 12GB of text.
(a) Starting interface of interactive engine, with first command typed, to stop the analysis at the first alarm.
(b) The engine jumped back to the beginning of the statement generating the alarm.
(c) Interval analysis: Inspecting abstract information related to __s.
(d) Relational analysis started, with breakpoints to reach the alarm detected in Figure 6(a).
(e) Relational analysis: Inspecting abstract information related to __s.
Abbildung 6: Interactive engine workflow example on the analysis of coreutils fmt.
In this section, we showcase an abstract debugger, providing an interactive interface to the user. The user decides how they want to navigate the abstract execution of the program, and computations are performed on-the-fly accordingly. This debugger provides an interface similar to gdb, except that it works on the abstract execution of the program while gdb would work on a concrete execution. As Mopsa currently works as an interpreter on the AST (and not as a generic equation solver), the flow of execution of the analysis is close to that of the concrete execution of the program, and easy to understand. The abstract execution can be navigated using diverse strategies, such as going to the next statement, entering inner analysis of functions, or continuing until the next breakpoint. It is also possible to observe intra-instruction analysis, and rewriting operations that are at the heart of Mopsa. Breakpoints can be program locations, functions, transfer functions or the next detected alarm. The abstract state can be printed, as well as projections of abstract information related to selected abstract variables.
Combined with a terminal multiplexer such as tmux, this interactive interface can be used to perform some side-by-side debugging of the analysis of a same program in two different configurations. Thanks to some wrapper scripts we have developed, commands for the interactive engine have to be entered once and will be given to both sides of the analysis, further easing debugging.
We show how the interactive engine can be used in Figure 6, on an example where we analyze coreutils fmt. The analysis is fully context-sensitive, and supposes symbolic arguments are passed to fmt.
The interactive engine is shown just after it was started in Figure 6(a). The prompt (represented by mopsa >>) has been given two commands: breakpoint #alarm and continue (abbreviated as b #a and c, and chained using a semicolon). This adds a breakpoint at the next encountered alarm and runs the analysis until this breakpoint is reached. The analysis raises an alarm in a call to builtin function strrchr at progname.c:59.
[…]
Let us restart the analysis in a relational setting, relying on the polyhedra abstract domain. In order to scale we use static packing (introduced by Bertrane et al. (2015)), a technique that keeps multiple polyhedra of small dimensions rather than a high-dimension polyhedra. We introduce here a specific pack handling the symbolic arguments used to analyze the program. In Figure 6(d), we add breakpoints to reach the program location where the alarm was raised with the interval analysis (Figure 6(a)).
If we break to the program position where the alarm was raised with the interval analysis, we notice the numeric relations of Figure 6(e) express exactly what is needed to prove the existentially quantified formula from Figure 6(b). This change of configuration means the first alarm has been removed by moving to a more expressive domain.
(a) The center panel shows source code annotated with one of the alarms reported by Mopsa through the LSP, and the list of alarms in the bottom panel.
(b) Using the DAP, a user can set a breakpoint at line 325 (right panel), and then inspect the program state (left panel).
Abbildung 7: Analysis of coreutils fmt through VSCode’s interface.
From command-line interfaces to IDEs.
The whole interface of Mopsa, from its batch mode providing alarm reports to its abstract debugger, are available through the command-line. We provide similar interfaces for IDE users, by leveraging the Language Server Protocol (LSP) and the Debug Adapter Protocol (DAP) originally developed by VSCode. In particular, we can report results of our analysis directly on the source code thanks to the LSP Figure 7(a), just as linters do. We use the DAP to provide interactive, abstract debugger sessions within the IDE, highlighted in Figure 7(b).
6 Automated Testcase Reduction
Testcase reduction is an automated approach aiming at minimizing a test while keeping a specific property. One tool performing automated testcase reduction for C is creduce, developed by Regehr et al. (2012). It has originally been applied to compilers (where a test is an input program), and enjoys widespread use in this case. For example, GCC Wiki Contributors (2011) provide a guide to testcase reduction and asks for reduced testcases in their bug reports. We report our successful use of creduce on static analyzers, which are similarly highly complex pieces of code, manipulating potentially large input programs.
The creduce workflow is summarized in Figure 8. The user provides an input testcase (file.c) exhibiting an unwanted behavior (such as a crash). An oracle, written as a shell script, describes whether the unwanted behavior still occurs in partially reduced versions. creduce will then loop until it reaches a minimal testcase satisfying the oracle. One challenge of putting automated testcase reduction into practice is to establish sufficiently robust, yet automated, testcase oracles.
We start by describing two ways to leverage automated testcase reduction to ease debugging in Mopsa in Section 6.1. Our first usecase concerns internal errors from the analyzer that can happen deep into long analyses. Our second usecase pinpoints soundness issues by comparing two analyses, yielding different results on a same program. We finish by highlighting in Section 6.2 that the collaboration between Mopsa and creduce is a two-way street, as Mopsa can simplify testcase reduction for multi-file projects with complex commands such as GNU coreutils.
file.coracle.shcreducesmall.c
Abbildung 8: creduce pipeline.
6.1 Leveraging Automated Testcase Reduction
Internal error reduction.
We have successfully used creduce to pinpoint and fix internal errors within Mopsa which result in uncaught OCaml exceptions terminating the analysis. In that case, the oracle we use first checks that the produced file can be compiled without errors666creduce can produce C programs with illegal syntax or types, that are naturally rejected (early) by a compiler’s frontend., and then checks that the same internal error is raised during the analysis of the reduced testcase. In our experience, this kind of reduction only requires the exception name and message to be passed to the oracle. We provide in Table 4 examples of issues or merge requests where automated testcase reduction has been used. These cases stem from benchmarks from the Software-Verification Competition, or from coreutils programs. Lines of code correspond to the number of lines of preprocessed C code (cf. Section 6.2), formatted using clang-format, and measured through command wc -l. In our experience, reduction time is less than 12 hours, meaning that it can easily be run overnight and save a considerable amount of human time. It helped us solve issues that were otherwise out of reach. For example, issue 81 from Table 4 was reduced from one linux device driver code from the Software Verification Competition, and caused by an implementation bug in the string length domain from Journault et al. (2018).
Reference Original LoC Reduced LoC Reduction
Issue 76 28,737 18 99.94%
Issue 81 15,627 8 99.95%
Issue 134 17,411 10 99.94%
Issue 135 7,016 12 99.83%
M.R. 130 77,981 20 99.97%
M.R. 145 77,427 19 99.98%
Tabelle 4: Internal error reduction examples, taken from the Gitlab repository of Mopsa.
Differential-configuration reduction.
The testcase reduction for internal errors is the canonical usecase of such tools. We have had some recent successes in differential-configuration reduction, easing the debugging of cases where two different configurations of Mopsa (Section 2.2) yield contradictory analysis results on a given program. We have typically applied it when an analysis is unsound, and where the culprit (abstract domain or reduction) is included in one configuration and not the other. This is particularly useful given the large number of combinations of abstract domains Mopsa enables through its modular design. We have used this approach to simplify some soundness issues reported by external users (# 179, # 182), who then integrated it in their process when reporting further issues (# 184, # 185). Thanks to this approach, we have minimal programs in which to debug the source of unsoundness. For example, the reduction from issue 182 allowed to quickly identify the division from the integer powerset abstraction as the source of unsoundness which was then easily fixed. We have currently not explored how creduce could help us pinpoint precision improvements.
6.2 Leveraging Mopsa to Ease Multi-file Reduction
One of the current usability barriers to automated testcase reduction through creduce and its sibling cvise (developed by Pflanzer and Liška (2024)) is the support of multi-file projects. Indeed, creduce requires the explicit list of files to be reduced [^7] https://github.com/csmith-project/creduce/blob/31e855e290970cba0286e5032971509c0e7c0a80/creduce/creduce.in#L197. This list can be difficult to establish on large open-source projects, such as coreutils, where a build system like make takes care of compiling the various sources into an executable, through a list of complex rules. In addition, some large projects may use different files with different compilation options, which would create an additional difficulty in using standard creduce.
Mopsa natively supports the analysis of multi-file C projects, through a utility called mopsa-build which creates a compilation database by instrumenting the compilation process. This process can include configure scripts, make, cmake, etc. mopsa-build overrides environment variables to record all compiler and linker calls, as well as the options that were passed to them. Due to its seamless nature, mopsa-build can be used as a drop-in replacement of various build systems such as make: mopsa-build make. Then, this compilation database can be leveraged by the C analysis to analyze a specific target of the build system. An important side-effect of this process is the ability of the analysis to generate a single preprocessed file [^8] The single file is generated as a kind of source-level (or AST-level) linking., which does heavily simplify automated testcase reduction.
7 Related Work
To the best of our knowledge Andreasen et al. (2017) are the first to describe approaches they use to improve their static analysis of JavaScript. They describe combinations of techniques relying on delta-debugging, soundness testing (through comparison with concrete values) and blended analysis (injection of concrete values to restrict the abstract state). The automated testcase reduction we use is close to the delta-debugging techniques they rely on, but otherwise our approaches seem complementary.
Static analyzer interfaces.
Our interactive engine can act as an abstract debugger, either from the command-line interface, or through IDEs supporting the debug-adapter protocol (DAP). Some other analyzers can report their alarms through the language-server protocol (LSP) developed initially for linters. A specific interface called MagpieBridge has been developed by Luo et al. (2019) to simplify the integration of static analyzers within LSP. Molle et al. (2023) showcase a cross-level debugger, working on the analyzed program and enabling conditional breakpoints expressed either on the analyzed program or the state of the analyzer itself. The Goblint static analyzer provides a different kind of abstract debugger developed by Holter et al. (2024). It provides an interactive, graphical exploration of the abstract states in the control-flow graph of the program once the analysis has successfully finished. Both abstract debuggers from Goblint and Mopsa provide an IDE integration through the Debug Adapter Protocol. In their documentation of the Goblint analyzer, Saan et al. (2024b) also mention automated testcase reduction, in particular to debug fixpoint termination issues.
A blog post from the Frama-C team by Maroneze (2020) highlights that automated testcase reduction simplified their interaction with industrial clients having private codebases. At least for runtime errors, these industrial clients can run automated testcase reduction by themselves. The reduction will yield a highly simplified testcase that will not leak important parts of the initial, private codebase, which can then be sent to Frama-C developers without any confidentiality issues.
Testing the soundness and precision of static analyzers.
Bugariu et al. (2018) perform automated testing of numerical abstract domains, by checking that some chosen properties should be verified. We rely on a similar yet simplified approach in Section 4.3, through an encoding of a heuristic rule to detect unsoundness during an analysis. However, our approach is not specific to numerical abstract domains.
Klinger et al. (2019) describe ways to automatically compare the soundness and precision of different C static analyzers on programs from the Software-Verification Competition (SV-Comp). In particular, the original program can be mutated to check the results at different program points, and a notion of $\delta-$unsoundness is established, to reduce spurious warnings. In our case, we have only considered different configurations of Mopsa, but not compared it using testcase reduction to other static analysis tools. Taneja et al. (2020) develop SMT-based algorithms that can detect soundness and precision errors of some of LLVM dataflow analyses.
Formally verified static analyzers, such as the work of Jourdan et al. (2015), will not require debugging of unsound results by construction. However, we believe the techniques we presented could still be interesting to investigate precision issues.
8 Conclusion
This article documents and shares the practices we have established for the maintenance of the Mopsa static analyzer during the last 7 years. In particular, we rely on a measure of precision than can be computed automatically on software without baselines on the number of true bugs. This approach increases the transparency of the analysis, and simplifies regression detection. We have shown different tools (profiler, debugger) focusing on the abstract execution of the analyzed program, and reported use of automated testcase reduction to simplify our debugging. Following the work of Andreasen et al. (2017), we hope this article will inspire other groups and encourage other researchers to document and share their practices.
There are still some challenges revolving around the development of Mopsa. The number of different configurations to analyze a given language can grow quite quickly due to the modular architecture of Mopsa. We are currently performing regression tests on selected, specific configurations to reduce the computational cost. Code maintenance and debugging can still take some time, and onboarding material takes a lot of time to create and maintain. Finally, we are looking into providing an install-free version of Mopsa – through a web page for example – meaning that prospective users can quickly test it without any installation.
Literatur
  • Andreasen et al. (2017) Andreasen ES, Møller A, Nielsen BB (2017) Systematic approaches for increasing soundness and precision of static analyzers. In: Ali K, Cifuentes C (eds) Proceedings of the 6th ACM SIGPLAN International Workshop on State Of the Art in Program Analysis, SOAP@PLDI 2017, Barcelona, Spain, June 18, 2017, ACM, pp 31–36, DOI: 10.1145/3088515.3088521
  • Arceri et al. (2023) Arceri V, Dolcetti G, Zaffanella E (2023) Speeding up static analysis with the split operator. In: Ferrara P, Hadarean L (eds) Proceedings of the 12th ACM SIGPLAN International Workshop on the State Of the Art in Program Analysis, SOAP 2023, Orlando, FL, USA, 17 June 2023, ACM, pp 14–19, DOI: 10.1145/3589250.3596141
  • Bau et al. (2022) Bau G, Miné A, Botbol V, Bouaziz M (2022) Abstract interpretation of Michelson smart-contracts. In: Proc.~of the 11th ACM SIGPLAN International Workshop on the State Of the Art in Program Analysis (SOAP’22), ACM, pp 36–43, DOI: 10.1145/3520313.3534660
  • Becchi and Zaffanella (2020) Becchi A, Zaffanella E (2020) Pplite: Zero-overhead encoding of NNC polyhedra. Inf Comput 275:104620, DOI: 10.1016/J.IC.2020.104620
  • Bertrane et al. (2015) Bertrane J, Cousot P, Cousot R, Feret J, Mauborgne L, Miné A, Rival X (2015) Static analysis and verification of aerospace software by abstract interpretation. Found Trends Program Lang 2(2-3):71–190, DOI: 10.1561/2500000002
  • Black (2018) Black PE (2018) Juliet 1.3 test suite: Changes from 1.2. US Department of Commerce, National Institute of Standards and Technology
  • Bugariu et al. (2018) Bugariu A, Wüstholz V, Christakis M, Müller P (2018) Automatically testing implementations of numerical abstract domains. In: Huchard M, Kästner C, Fraser G (eds) Proceedings of the 33rd ACM/IEEE International Conference on Automated Software Engineering, ASE 2018, Montpellier, France, September 3-7, 2018, ACM, pp 768–778, DOI: 10.1145/3238147.3240464
  • Bühler et al. (2024) Bühler D, Maroneze A, Perrelle V (2024) Abstract Interpretation with the EvaEva (Frama-C plug-in) Plug-in, Springer International Publishing, Cham, pp 131–186. DOI: 10.1007/978-3-031-55608-1_3
  • Cousot and Cousot (1977) Cousot P, Cousot R (1977) Abstract interpretation: A unified lattice model for static analysis of programs by construction or approximation of fixpoints. In: POPL, ACM, pp 238–252, DOI: 10.1145/512950.512973
  • Danial (2021) Danial A (2021) cloc: v1.92. DOI: 10.5281/zenodo.5760077
  • Delmas et al. (2021) Delmas D, Ouadjaout A, Miné A (2021) Static analysis of endian portability by abstract interpretation. In: SAS, Springer, Lecture Notes in Computer Science, vol 12913, pp 102–123, DOI: 10.1007/978-3-030-88806-0_5
  • GCC Wiki Contributors (2011) GCC Wiki Contributors (2011) Gcc wiki – a guide to testcase reduction. URL https://gcc.gnu.org/wiki/A_guide_to_testcase_reduction
  • Gregg (2016) Gregg B (2016) The flame graph. Commun ACM 59(6):48–57, DOI: 10.1145/2909476, URL https://doi.org/10.1145/2909476
  • Gurfinkel and Navas (2021) Gurfinkel A, Navas JA (2021) Abstract interpretation of LLVM with a region-based memory model. In: VSTTE, Springer, Lecture Notes in Computer Science, vol 13124, pp 122–144, DOI: 10.1007/978-3-030-95561-8_8
  • Gurfinkel et al. (2015) Gurfinkel A, Kahsai T, Komuravelli A, Navas JA (2015) The seahorn verification framework. In: CAV (1), Springer, Lecture Notes in Computer Science, vol 9206, pp 343–361, DOI: 10.1007/978-3-319-21690-4_20
  • Helm et al. (2024) Helm D, Keidel S, Kampkötter A, Düsing J, Roth T, Hermann B, Mezini M (2024) Total recall? how good are static call graphs really? In: ISSTA 2024
  • Holter et al. (2024) Holter K, Hennoste JO, Lam P, Saan S, Vojdani V (2024) Abstract debuggers: Exploring program behaviors using static analysis results. In: Edwards J (ed) Proceedings of the 2024 ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software, Association for Computing Machinery, Onward! 2024, to appear
  • Jeannet and Miné (2009) Jeannet B, Miné A (2009) Apron: A library of numerical abstract domains for static analysis. In: CAV, Springer, Lecture Notes in Computer Science, vol 5643, pp 661–667, DOI: 10.1007/978-3-642-02658-4_52
  • Jourdan et al. (2015) Jourdan J, Laporte V, Blazy S, Leroy X, Pichardie D (2015) A formally-verified C static analyzer. In: POPL, ACM, pp 247–259, DOI: 10.1145/2676726.2676966
  • Journault et al. (2018) Journault M, Miné A, Ouadjaout A (2018) Modular static analysis of string manipulations in C programs. In: Podelski A (ed) Static Analysis - 25th International Symposium, SAS 2018, Freiburg, Germany, August 29-31, 2018, Proceedings, Springer, Lecture Notes in Computer Science, vol 11002, pp 243–262, DOI: 10.1007/978-3-319-99725-4_16
  • Journault et al. (2019) Journault M, Miné A, Monat R, Ouadjaout A (2019) Combinations of reusable abstract domains for a multilingual static analyzer. In: VSTTE, Springer, Lecture Notes in Computer Science, vol 12031, pp 1–18, DOI: 10.1007/978-3-030-41600-3_1
  • Klinger et al. (2019) Klinger C, Christakis M, Wüstholz V (2019) Differentially testing soundness and precision of program analyzers. In: Zhang D, Møller A (eds) Proceedings of the 28th ACM SIGSOFT International Symposium on Software Testing and Analysis, ISSTA 2019, Beijing, China, July 15-19, 2019, ACM, pp 239–250, DOI: 10.1145/2644805
  • Livshits et al. (2015) Livshits B, Sridharan M, Smaragdakis Y, Lhoták O, Amaral JN, Chang BE, Guyer SZ, Khedker UP, Møller A, Vardoulakis D (2015) In defense of soundiness: a manifesto. Commun ACM 58(2):44–46, DOI: 10.1145/2644805
  • Luo et al. (2019) Luo L, Dolby J, Bodden E (2019) Magpiebridge: A general approach to integrating static analyses into ides and editors (tool insights paper). In: Donaldson AF (ed) 33rd European Conference on Object-Oriented Programming, ECOOP 2019, July 15-19, 2019, London, United Kingdom, Schloss Dagstuhl - Leibniz-Zentrum für Informatik, LIPIcs, vol 134, pp 21:1–21:25, DOI: 10.4230/LIPICS.ECOOP.2019.21
  • Maroneze (2020) Maroneze A (2020) Debugging Frama-C analyses: Better privacy with C-reduce. URL https://frama-c.com/2020/04/02/creduce.html
  • Merigoux et al. (2021) Merigoux D, Chataing N, Protzenko J (2021) Catala: a programming language for the law. Proc ACM Program Lang 5(ICFP):1–29, DOI: 10.1145/3473582
  • Milanese and Miné (2024) Milanese M, Miné A (2024) Generation of violation witnesses by under-approximating abstract interpretation. In: VMCAI (1), Springer, Lecture Notes in Computer Science, vol 14499, pp 50–73, DOI: 10.1007/978-3-031-50524-9_3
  • Miné et al. (2024) Miné A, Ouadjaout A, Journault M, Fromherz A, Monat R, Parolini F, Milanese M, Boillot J (2024) Mopsa: Modular open static analysis platform. URL https://gitlab.com/mopsa/mopsa-analyzer/
  • Molle et al. (2023) Molle MV, Vandenbogaerde B, Roover CD (2023) Cross-level debugging for static analysers. In: Saraiva J, Degueule T, Scott E (eds) Proceedings of the 16th ACM SIGPLAN International Conference on Software Language Engineering, SLE 2023, Cascais, Portugal, October 23-24, 2023, ACM, pp 138–148, DOI: 10.1145/3623476.3623512
  • Monat (2021) Monat R (2021) Static type and value analysis by abstract interpretation of python programs with native C libraries. (analyse statique, de type et de valeur, par interprétation abstraite, de programmes python utilisant des librairies C). PhD thesis, Sorbonne University, Paris, France, URL https://tel.archives-ouvertes.fr/tel-03533030
  • Monat et al. (2020a) Monat R, Ouadjaout A, Miné A (2020a) Static type analysis by abstract interpretation of python programs. In: ECOOP, Schloss Dagstuhl - Leibniz-Zentrum für Informatik, LIPIcs, vol 166, pp 17:1–17:29, DOI: 10.4230/LIPICS.ECOOP.2020.17
  • Monat et al. (2020b) Monat R, Ouadjaout A, Miné A (2020b) Value and allocation sensitivity in static python analyses. In: SOAP@PLDI, ACM, pp 8–13, DOI: 10.1145/3394451.3397205
  • Monat et al. (2021) Monat R, Ouadjaout A, Miné A (2021) A multilanguage static analysis of python programs with native C extensions. In: SAS, Springer, Lecture Notes in Computer Science, vol 12913, pp 323–345, DOI: 10.1007/978-3-030-88806-0_16
  • Monat et al. (2023) Monat R, Ouadjaout A, Miné A (2023) Mopsa-c: Modular domains and relational abstract interpretation for C programs (competition contribution). In: TACAS Part II, Springer, Lecture Notes in Computer Science, vol 13994, pp 565–570, DOI: 10.1007/978-3-031-30820-8_37
  • Monat et al. (2024a) Monat R, Fromherz A, Merigoux D (2024a) Formalizing date arithmetic and statically detecting ambiguities for the law. In: ESOP (2), Springer, Lecture Notes in Computer Science, vol 14577, pp 421–450, DOI: 10.1007/978-3-031-57267-8_16
  • Monat et al. (2024b) Monat R, Milanese M, Parolini F, Boillot J, Ouadjaout A, Miné A (2024b) Mopsa-c: Improved verification for C programs, simple validation of correctness witnesses (competition contribution). In: TACAS Part III, Springer, Lecture Notes in Computer Science, vol 14572, pp 387–392, DOI: 10.1007/978-3-031-57256-2_26
  • Namjoshi and Pavlinovic (2018) Namjoshi KS, Pavlinovic Z (2018) The impact of program transformations on static program analysis. In: SAS, Springer, Lecture Notes in Computer Science, vol 11002, pp 306–325, DOI: 10.1007/978-3-319-99725-4_19
  • Ouadjaout and Miné (2020) Ouadjaout A, Miné A (2020) A library modeling language for the static analysis of C programs. In: Pichardie D, Sighireanu M (eds) SAS, Springer, Lecture Notes in Computer Science, vol 12389, pp 223–247, DOI: 10.1007/978-3-030-65474-0_11
  • Parolini and Miné (2024) Parolini F, Miné A (2024) Sound abstract nonexploitability analysis. In: VMCAI (2), Springer, Lecture Notes in Computer Science, vol 14500, pp 314–337, DOI: 10.1007/978-3-031-50521-8_15
  • Pflanzer and Liška (2024) Pflanzer M, Liška M (2024) CVise: Super-parallel python port of the C-Reduce. URL https://github.com/marxin/cvise/
  • Regehr et al. (2012) Regehr J, Chen Y, Cuoq P, Eide E, Ellison C, Yang X (2012) Test-case reduction for C compiler bugs. In: Vitek J, Lin H, Tip F (eds) ACM SIGPLAN Conference on Programming Language Design and Implementation, PLDI ’12, Beijing, China - June 11 - 16, 2012, ACM, pp 335–346, DOI: 10.1145/2254064.2254104
  • Saan et al. (2024a) Saan S, Schwarz M, Erhard J, Tilscher S, Holter K, Vogler R, Apinis K, Vojdani V (2024a) Goblint. DOI: 10.5281/zenodo.5735006, URL https://github.com/goblint/analyzer
  • Saan et al. (2024b) Saan S, Schwarz M, Erhard J, Tilscher S, Holter K, Vogler R, Apinis K, Vojdani V (2024b) Goblint analyzer documentation on testcase reduction. URL https://goblint.readthedocs.io/en/stable/developer-guide/debugging/
    #debugging-issues-with-larger-programs
  • Smaragdakis et al. (2011) Smaragdakis Y, Bravenboer M, Lhoták O (2011) Pick your contexts well: understanding object-sensitivity. In: POPL, ACM, pp 17–30, DOI: 10.1145/1926385.1926390
  • Sotin (2010) Sotin P (2010) Quantifying the Precision of Numerical Abstract Domains. Research report, URL https://inria.hal.science/inria-00457324
  • Taneja et al. (2020) Taneja J, Liu Z, Regehr J (2020) Testing static analyses for precision and soundness. In: CGO ’20: 18th ACM/IEEE International Symposium on Code Generation and Optimization, San Diego, CA, USA, February, 2020, ACM, pp 81–93, DOI: 10.1145/3368826.3377927
[…]